OpenStudy (anonymous):

HELP ME!!! find the length of the chord common to the parabolas y^2 = 2x + 4y + 6 and y^2 = 3x + 3y +1

OpenStudy (mandre):

I drew your graphs in desmos.com and they don't seem to have a common chord, unless I misunderstand your question?

OpenStudy (anonymous):

it is really written on the book. hehe

OpenStudy (mandre):

Don't mind me, I'm confused lol.

OpenStudy (anonymous):

you may isolate x in both equations and find intersect points with equating x's in terms of y

OpenStudy (dumbcow):

find the 2 intersection points of 2 parabolas, then use distance formula to get length of chord

OpenStudy (anonymous):

thank you :)

OpenStudy (campbell_st):

actually the 2 parabola's intersect the easy way to find the equation of the chord if to equate the 2 equations are they are both equal to y^2 then simplify and you'll get the equation of the chord

OpenStudy (campbell_st):

you don't need to find points of intersection... or anything else... is a difficult problem to understand by simple to get the solution

OpenStudy (anonymous):

can you show it to me step by step?

OpenStudy (campbell_st):

ok so the equations are \[y^2 = 2x + 4y + 6... and..... y^2 = 3x + 3y + 1\] so you end up with \[2x + 4y + 6 = 3x + 3y + 1\] its that simple

OpenStudy (anonymous):

\[10\sqrt{2}\]

OpenStudy (anonymous):

is the answer

OpenStudy (campbell_st):

here is the graph

OpenStudy (campbell_st):

ops I read equation rather than length...sorry about that

OpenStudy (campbell_st):

so just find where the chord intersects... and you'll get the 2 points for the distance

OpenStudy (anonymous):

@zith how did you get that 10 square root blabla? :)

OpenStudy (campbell_st):

so you are looking at \[(x -5)^2 = 3x + 3(x -5) + 1\] just work on that...

OpenStudy (anonymous):

@campbell_st how?

OpenStudy (mandre):

Simplify campbell's last equation into the form y = mx + b then use that equation and the equation of one of the parabolas to solve for x and y. You should end up with 2 points (x, y). Use those points and pythagoras to calculate the length.

OpenStudy (campbell_st):

\[x^2 -10x + 25 = 3x + 3x -15 + 1... or .... x^2 -16x + 39 = 0\] which factors to (x - 3)(x - 13) = 0 so the x values are x = 3 and x = 13 so the y values are y = -2 and y = 8 so the points are (3, -2) and (13, 8) find the distance between them

OpenStudy (anonymous):

so i have already told u about distance formula it is

OpenStudy (anonymous):

Thank you so much @Mandre :)

OpenStudy (anonymous):